# LVLM-COUNT: Enhancing the Counting Ability of Large Vision-Language Models

### v1

@@ 3.4 Target Counting
At this point, all the sub-images obtained from the cropped areas are gathered. Since these are small partitions of the original image, they might lack the desired resolution. Thus, using a super-resolution model can alleviate this problem for the sub-images. In this work, we use Real-ESRGAN (Wang et al., 2021). Then, question $Q$ and each sub-image are given as input to an LVLM. At the end of the loop, the recorded numbers for the sub-images are aggregated to form the final answer.  [^1] Note that using super-resolution before counting the objects in the sub-images is also the default behavior of our method. For a specific dataset where we observe that the resolution is preserved after division, we deactivate super-resolution, as the cost and usage of LVLM APIs is much higher. Finally, for images with a very large number of objects, sometimes LVLMs refuse to count, citing the large number. In those cases, an “estimate” prompt is given instead of a “count” prompt..
@@ 4 Experiments
In this section, we present the performance results of our method on a counting-specific dataset, an open-ended counting benchmark with two types of questions, simple and complex, and a challenging counting benchmark that we propose using emoji icons. We compare the results to state-of-the-art models which have been specifically trained on counting datasets, and we also compare to state-of-the-art models which have not been trained on counting datasets. The code to reproduce the experiments can be found here: https://github.com/mrghofrani/lvlm-count.
@@ 4.1 Datasets and benchmarks
FSC-147 (Ranjan et al., 2021). FSC-147 is a counting dataset that contains $6135$ images, spanning $147$ different object categories such as kitchen utensils, office supplies, vehicles, and animals. The number of objects in each image ranges from $7$ to $3731$, with an average of $56$ objects per image. The dataset is split into training, validation, and test sets. A total of $89$ object categories are assigned to the training set, $29$ to the validation set, and $29$ to the test set, with different categories in each split. The training set contains $3659$ images, with the validation and test sets containing $1286$ and $1190$ images, respectively. For each image in the test set, a single category name is given, and the expected output is the number of instances from the category.
Open-ended Counting Benchmark. TallyQA (Acharya et al., 2019) is an open-ended counting dataset that includes complex counting questions involving relationships between objects, attribute identification, reasoning, and more. TallyQA is quite a large dataset, with the train set having $249,318$ questions and the test set having $22,991$ simple and $22,991$ complex counting questions. Please refer to Appendix C and Figure 9 for more information on simple and complex categorization of the questions. The number of objects in each image ranges from 00 to $15$. The dataset has a heavy bias towards a low number of objects (see Figure 8 in the Appendix). To alleviate this bias and create a benchmark for efficiently measuring the simple and complex open-ended capabilities of a counting model, we randomly sample $10$ questions per ground truth count. This results in $155$ simple and $149$ complex open-ended counting questions in total. It is important to note that the bias in TallyQA is pronounced; for most ground truth values greater than $10$, there are fewer than $10$ samples available in the entire test set.
Emoji-Count. Although TallyQA addresses the scarcity of complex counting questions in prior datasets to a certain degree, the range of target objects is limited, spanning from 00 to $15$. To our knowledge, no counting benchmark exists for large numbers involving complex reasoning. To this end, we propose a challenging counting benchmark using emoji icons. From the $1816$ standard emoji icons, we remove those that directly overlap with concepts demonstrated by other icons. We then group the remaining $1197$ icons into $82$ classes. In each class, there are icons from the same or similar object categories, but with subtle differences that require complex reasoning to distinguish. For each of the $82$ classes, an empty $1024\times 1024$ image is first created. This image is filled with six categories chosen randomly from the class, with each category having a random count between $30$ and $50$ in the image. We illustrate an example of this dataset in Figure 7.
(a) Q: How many waning gibbous moons are there in the image? Answer: 13.
@@
Figure 7: Illustration of two challenging cases from Emoji-Count. Note that, for convenience in visualization, a smaller version of the real images is depicted here. The real images are $1024\times 1024$ and contain many more icons. In Figure 7(a), the class name is “Moon Phase”. To predict the correct answers for the question, the model needs accurate information about the different moon phases in addition to strong counting capability. In Figure 7(b), the class name is “Clock Time”. In this case, the model needs to be able to read the time shown on the clocks in addition to strong counting capability.
@@ 4.2 Results
The following discusses the numerical results of our experiments with LVLM-Count on each benchmark described in Section 4.1. For visual examples of LVLM-Count’s performance on each benchmark, see Appendix M. Additionally, for the ablation study and experiments on the PASCAL VOC dataset (Everingham et al., 2015), see Appendix A and Appendix K, respectively.
FSC-147. We compare the performance of LVLM-Count to an extensive list of state-of-the-art counting models on the test set of the FSC-147 dataset. We also include two baselines: i) taking the number of target segmentation masks as the final answer, and ii) giving the output of the target segmentation stage to base GPT-4o and asking it to count the masks. We run different experiments using GPT-4o, Gemini 1.5 Pro, and an open-source model Qwen2 VL 72B AWQ (Yang et al., 2024) as LVLMs. Since objects in the FSC-147 dataset are uniformly distributed across scenes, we deactivate the cluster-based method for finding division points and instead use two vertical and two horizontal object-aware division paths with equidistant start and end points. We also deactivate the super-resolution. We further run an experiment with GPT-4o as the LVLM that uses the cluster-based approach along both horizontal and vertical axes to automatically determine the start and end points of the division paths. This is done to measure the performance of our method when prior knowledge is not used to set the start and end points. The rest of the process follows the procedure described in Section 3. The expression $E$ used in different stages of our method is the category name provided in the test set. A simple $Q$ in the form of “How many $E$ are there? If you don’t see any, say zero.” is built and given as a text prompt to the LVLM during the counting stage. The results are shown in Table 1. For this dataset, the mean absolute error (MAE) and root mean square error (RMSE) are reported. We observe that our method enhances the performance of all three LVLMs in terms of MAE. With our method, all three LVLMs outperform TFOC, which which is a training-free method. Interestingly, although the base Qwen2 VL 72B AWQ is not as powerful as its commercial counterpart, GPT-4o, it performs almost on par with GPT-4o when both use our pipeline. The results also show that even without leveraging prior information to set the start and end points of the division paths, our pipeline remains effective.
@@
Table 1: Evaluation of state-of-the-art models on the test set of the FSC-147 dataset. The column “Trained Model” indicates if a model has been trained on FSC-147. In all tables, the results for base LVLMs and LVLM-Count are reported over three trials. Columns marked with $\Delta$ show the improvement that LVLM-Count brings over the base LVLM that it uses. For accuracy metrics, refer to Table 9. Additionally, MAE analysis across different intervals of ground truth values is provided in Appendix H.
Method Trained Model MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC (Shi et al., 2024) 24.79 - 137.15 - VLCounter (Kang et al., 2024) 17.05 - 106.16 - CounTX(Amini-Naieni et al., 2023) 15.88 - 106.29 - $\text{DAVE}_{\text{prm}}$ (Pelhan et al., 2024) 14.90 - 103.42 - CountGD (Amini-Naieni et al., 2024) 14.76 - 120.42 - GroundingREC (Dai et al., 2024) 10.12 - 107.19 - Number of target segmentaion masks 44.14 - 154.39 - Base GPT-4o counting the target segmentaion masks 38.45 - 38.45 - GPT-4o 23.75 - 137.39 - LVLM-Count (GPT-4o as LVLM, Cluster-based on both axes) 17.67 $\downarrow 6.08$ 90.61 $\downarrow 46.78$ LVLM-Count (GPT-4o as LVLM) 16.23 $\downarrow 7.52$ 76.09 $\downarrow 61.3$ Gemini 1.5 Pro 25.20 - 108.76 - LVLM-Count (Gemini 1.5 Pro as LVLM) 14.85 $\downarrow 10.35$ 85.60 $\downarrow 23.16$ Qwen2 VL 72B AWQ 33.45 - 145.90 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 17.77 $\downarrow 15.68$ 113.06 $\downarrow 32.84$
Table 1 shows that, although we outperform models that have not been trained on FSC-147, the best-performing models are those that have been trained on this dataset. However, in subsequent experiments, we demonstrate that, while the best-performing models on FSC-147 show acceptable generalization on benchmarks they have not been trained on, they lack the generalization ability of our method on new datasets, and they are even outperformed by the base LVLM models.
Open-ended Counting Benchmark. We evaluate the performance of LVLM-Count on the simple and complex questions in the open-ended counting benchmark mentioned above. For LVLM-Count, we compute the start and end points of the object-aware division paths using the unsupervised and parameter-free clustering approach described in Section 3.3 on $x$-axis. This approach is necessary because the images in this benchmark are diverse, and objects are not uniformly distributed across the image. Thus, using a fixed number of object-aware division paths with equidistant start and end points does not perform well. We also conduct an experiment using GPT-4o as the LVLM, where we use the clustering method along both axes to find the start and end points.
Since these are special cases of VQA tasks, we report exact accuracy (EA), which is the standard performance metric for VQA, in addition to MAE and RMSE. We compare our method against the base GPT-4o, Gemini 1.5 Pro, and Qwen2 VL 72B models, as well as three of the best-performing trained models from Table 1, namely GroundingREC, CountGD, and $\text{DAVE}_{\text{prm}}$ [^2] GroundingREC in Table 2 does not use the same model weights as those trained for FSC-147. It was trained on the REC-8K dataset (Dai et al., 2024), which is specifically designed for referring expression counting, making it stronger for complex counting questions. CountGD and $\text{DAVE}_{\text{prm}}$ use the same model weights as those used for FSC-147, as the authors have not provided model weights trained on a dataset with referring expressions., and the only prior training-free model, TFOC. Note that due to extremely poor performance, we do not provide the original question to these models. Instead, we assist them with the $E$ extracted in our pipeline. The results are shown in Table 2. Our method improves both MAE and EA over the base GPT-4o model on both complex and simple benchmarks.
In general, however, there is barely any improvement over base LVLMs on simple questions. This is because the questions are straightforward, and the ground truths are within [0, 15]. Based on our observations from Figure 2(c) and Figure 18, we expect the base LVLMs to perform better in such cases. However, for complex questions, there is consistent improvement in EA across all three LVLMs. Interestingly, despite the considerable gap between the EA of the base Qwen2 VL 72B AWQ and base GPT-4o on complex questions, using LVLM-Count enables this open-source model to achieve a higher EA than the base GPT-4o. Moreover, observe that GroundingREC, CountGD, and $\text{DAVE}_{\text{prm}}$ are outperformed by the LVLM-based models on both simple and complex categories. This is because they have not been specifically trained on TallyQA. The performance gap is more pronounced on complex questions.
Table 2: Evaluation of models on the TallyQA benchmark. Numbers in parentheses under the $\Delta$ signs show the performance difference between LVLM-Count and the base LVLM it uses. Green indicates improvement, while red represents degradation. For more details about the simple and complex types of questions, please refer to Appendix C and Figure 9. For additional accuracy measures, refer to Table 10 and 11.
Method Simple Questions Complex Questions EA (%$\uparrow$($\Delta$) MAE $\downarrow$($\Delta$) RMSE $\downarrow$($\Delta$) EA (%$\uparrow$($\Delta$) MAE $\downarrow$($\Delta$) RMSE $\downarrow$($\Delta$) TFOC (Shi et al., 2024) 6.45 (-) 8.28 (-) 14.79 (-) 1.34 (-) 12.41 (-) 22.80 (-) $\text{DAVE}_{\text{prm}}$ (Pelhan et al., 2024) 5.81 (-) 16.06 (-) 35.26 (-) 3.36 (-) 28.36 (-) 57.56 (-) GroundingREC (Dai et al., 2024) 23.87 (-) 2.90 (-) 4.71 (-) 16.78 (-) 5.83 (-) 10.13 (-) CountGD (Amini-Naieni et al., 2024) 41.94 (-) 2.37 (-) 4.56 (-) 5.37 (-) 9.78 (-) 17.21 (-) Number of the target segmentation masks 27.31 (-) 2.60 (-) 4.10 (-) 15.44 (-) 4.25 (-) 6.78 (-) Base GPT-4o counting the target segmentation masks 25.38 (-) 2.82 (-) 2.82 (-) 11.41 (-) 4.99 (-) 4.99 (-) GPT-4o 44.30 (-) 1.38 (-) 2.35 (-) 29.08 (-) 2.60 (-) 4.74 (-) LVLM-Count (GPT-4o as LVLM, Cluster-based on both axes) 41.29 ($\downarrow 3.01$) 1.82 ($\uparrow 0.44$) 3.33 ($\uparrow 0.98$) 28.41 ($\downarrow 0.67$) 3.18 ($\uparrow 0.58$) 8.91 ($\uparrow 4.17$) LVLM-Count (GPT-4o as LVLM, Cluster-based on $x$-axis) 44.73 ($\uparrow 0.43$) 1.18 ($\downarrow 0.20$) 2.06 ($\downarrow 0.29$) 34.68 ($\uparrow 5.6$) 2.28 ($\downarrow 0.32$) 4.18 ($\downarrow 0.56$) Gemini 1.5 Pro 47.10 (-) 1.08 (-) 1.87 (-) 25.73 (-) 2.13 (-) 3.36 (-) LVLM-Count (Gemini 1.5 Pro as LVLM, Cluster-based on $x$-axis) 45.16 ($\downarrow 1.94$) 1.43 ($\uparrow 0.35$) 4.72 ($\uparrow 2.85$) 26.62 ($\uparrow 0.89$) 2.79 ($\uparrow 0.66$) 4.70 ($\uparrow 1.34$) Qwen2 VL 72B AWQ 49.03 (-) 1.44 (-) 2.74 (-) 24.61 (-) 3.21 (-) 5.35 (-) LVLM-Count (Qwen2 VL 72B AWQ as LVLM, Cluster-based on $x$-axis) 41.72 ($\downarrow 7.31$) 1.66 ($\uparrow 0.22$) 3.41 ($\uparrow 0.67$) 30.65 ($\uparrow 6.04$) 2.47 ($\downarrow 0.74$) 4.35 ($\downarrow 1$)
Emoji-Count. We evaluate the performance of LVLM-Count on the Emoji-Count benchmark. The results are shown in Table 3. For LVLM-Count, we deactivate the cluster-based division points and, similar to the FSC-147 experiment, use two vertical and horizontal object-aware division paths with equidistant start and end points. We also run an experiment with GPT-4o where LVLM-Count uses the clustering approach along both axes to determine the start and end points of the division paths. We report MAE and RMSE and compare against the base LVLMs, TFOC, $\text{DAVE}_{\text{prm}}$, CountGD, and GroundingREC. Note that none of the models have been exposed to this benchmark. Furthermore, it is a challenging benchmark as it requires understanding complex concepts. We observe that prior counting models perform poorly because, for any object of interest in the image, these models tend to count all the objects and cannot distinguish between different icons. Nonetheless, the base LVLMs show reasonable performance, but since the number of objects of interest is large in this dataset, all three base LVLMs are outperformed by LVLM-Count. The results for the clustering approach show that LVLM-Count is quite effective even when no prior knowledge is used to set the start and end points of the division paths.
Table 3: Evaluation of state-of-the-art models on the Emoji-Count benchmark. Columns marked with $\Delta$ show the improvement LVLM-Count brings over the base LVLM it uses. Please refer to Table 12 for more accuracy metrics.
@@
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC (Shi et al., 2024) 64.64 - 87.45 - $\text{DAVE}_{\text{prm}}$ (Pelhan et al., 2024) 198.99 - 208.08 - CountGD (Amini-Naieni et al., 2024) 137.93 - 156.80 - GroundingREC (Dai et al., 2024) 36.16 - 51.88 - Number of the target segmentation masks 82.47 - 107.98 - Base GPT-4o counting the target segmentation masks 107.72 - 162.12 - GPT-4o 22.51 - 35.94 - LVLM-Count (GPT-4o as LVLM, Cluster-based on both axes) 11.10 $\downarrow 11.41$ 24.23 $\downarrow 11.71$ LVLM-Count (GPT-4o as LVLM) 9.82 $\downarrow 12.69$ 19.46 $\downarrow 16.48$ Gemini 1.5 Pro 18.17 - 27.83 - LVLM-Count (Gemini 1.5 Pro as LVLM) 14.35 $\downarrow 3.82$ 23.42 $\downarrow 4.41$ Qwen2 VL 72B AWQ 82.41 - 186.32 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 20.67 $\downarrow 61.73$ 34.48 $\downarrow 151.84$
5 Limitations and Future Work
A limitation of LVLM-Count is the area detection stage. If the cropped areas do not provide enough context, performance may suffer. This is especially true for complex questions that require understanding the relationships between all the objects and the background in an image. This opens up opportunities for future work on improved area detection methods or the use of a context provider to complement the area detection stage. Another limitation arises with images containing thousands of objects. After one iteration of division, a significant number of objects may still remain in the sub-images. A potential solution is to apply additional iterations of division on the sub-images; however, the low resolution of these sub-images may make this infeasible. Developing a solution to maintain resolution is another direction for future work. Finally, in some cases, sub-images do not contain any objects of interest. The LVLM occasionally predicts a non-zero value in such instances. This is a weakness of LVLMs that requires special consideration to improve their performance.

### v3

@@ 3.4 Target Counting
All the sub-images obtained from the cropped areas are gathered. Then, question $Q$ and each sub-image are given as input to an LVLM. At the end of the loop, the recorded numbers for the sub-images are aggregated to form the final answer. For images with a very large number of objects, sometimes LVLMs refuse to count, citing the large number. In those cases, a new prompt requires the model to give the closest estimate of the number.
@@ 4 Experiments
In this section, we present the performance results of different LVLMs and their enhanced performance results using our method on a counting-specific dataset, a counting benchmark taken from a popular vision datasets, and a challenging counting benchmark that we propose using emoji icons. Additionally, we show the success of our design in enhancing counting performance on one of the most challenging counting datasets that features heavy occlusions and complex backgrounds. The code to reproduce the experiments can be found here: https://github.com/mrghofrani/lvlm-count.
@@ 4.1 Datasets and Benchmarks
FSC-147 (Ranjan et al., 2021). FSC-147 is a counting dataset that contains $6135$ images, spanning $147$ different object categories such as kitchen utensils, office supplies, vehicles, and animals. The number of objects in each image ranges from $7$ to $3731$, with an average of $56$ objects per image. The dataset is split into training, validation, and test sets. A total of $89$ object categories are assigned to the training set, $29$ to the validation set, and $29$ to the test set, with different categories in each split. The training set contains $3659$ images, with the validation and test sets containing $1286$ and $1190$ images, respectively. For each image in the test set, a single category name is given, and the expected output is the number of instances.
PASCAL VOC Benchmark We build a counting benchmark from PASCAL VOC dataset (Everingham et al., 2015). Similar to Chattopadhyay et al. (2017), we choose PASCAL VOC 2007 among other variants. This variant contains a training set of 2501 images, a validation set of 2510 images, and a test set of 4952 images, with 20 object categories that remain consistent across the splits. Each image includes annotations for instances of the 20 object categories in the dataset. We first create 20 simple counting questions asking for the number of objects from each of the 20 categories for every image in the test set. Then, we randomly sample five questions from each ground truth count. This process resulted in 102 questions in total. Finally, we manually checked the ground truth counts and corrected them if required.
Emoji-Count. To our knowledge, no counting benchmark exists with large number of objects in a scene involving complex reasoning. To this end, we propose a challenging counting benchmark using emoji icons. From the $1816$ standard emoji icons, we remove those that directly overlap with concepts demonstrated by other icons. We then group the remaining $1197$ icons into $82$ classes. In each class, there are icons from the same or similar object categories, but with subtle differences that require complex reasoning to distinguish. For each of the $82$ classes, an empty $1024\times 1024$ image is first created. This image is filled with up to six categories chosen randomly from the class, with each category having a random count between $30$ and $50$ in the image. For each image, we create questions that ask the number of instances of the available categories in the image. This results in 415 image-question pairs. We illustrate two examples of this dataset in Figure 7.
(a) Q: How many waning gibbous moons are there in the image? Answer: 18.
@@
Figure 7: Illustration of a smaller version of two challenging cases from Emoji-Count. In Figure 7(a), the class name is “Moon Phase”. In Figure 7(b), the class name is “Clock Time”.
Penguin Benchmark. The challenging Penguin dataset Penguin Research (2016) consistently exhibits heavy occlusion and complex background patterns that can easily be mistaken for penguins Arteta et al. (2016). The test set of this dataset is quite large, containing thousands of images with penguin counts ranging from 0 to 213. To build a manageable benchmark, we randomly sample 100 images while preserving a balanced ground truth range. We refer the reader to Appendix D for more details about this benchmark.
@@ 4.2 Results
The following section presents the numerical results of our experiments with the base LVLMs and their corresponding LVLM-Count on each benchmark described in Section 4.1. For visual examples of LVLM-Count’s performance on these benchmarks, see Appendix B. The ablation study results are discussed in Appendix A. Additionally, Appendix K provides the performance of SOTA counting models on these benchmarks as a reference for evaluating the LVLMs’ performance. Also, an inference time analysis is provided in Appendix J.
FSC-147. We evaluate the performance of various LVLMs and their enhanced performance using our method, LVLM-Count, on the test set of the FSC-147 dataset. We consider two baselines: (i) the baseline where the number of target segmentation masks is taken as the final answer, and (ii) the baseline where an image of the generated segmentation masks is provided to GPT-4o with a prompt to count the number of masks. Our experiments involve GPT-4o (a leading proprietary model), as well as two open-source models: Qwen2 VL 72B AWQ (Yang et al., 2024) and Gemma 3 27B Team et al. (2025). The expression $E$ used in different stages of our method corresponds to the category name provided in the test set. A simple query $Q$ in the form of “How many $E$ are there?” is constructed and used as the text prompt for the LVLM during the counting stage. In all experiments, the detection thresholds are set to $0.1$. The results are presented in Table 1, reporting the mean absolute error (MAE) and root mean square error (RMSE). Our findings indicate that LVLM-Count improves the performance of all three LVLMs in terms of MAE. Interestingly, while the base Qwen2 and Gemma 3 models are much less powerful than their commercial counterpart, they outperform the base GPT-4o when integrated into our pipeline.
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ Number of target segmentaion masks 44.14 - 154.39 - GPT-4o counting target segmentaion masks 38.45 - 156.11 - GPT-4o 25.17 - 137.86 - LVLM-Count (GPT-4o as LVLM) 18.09 $\downarrow 7.08$ 88.77 $\downarrow 49.09$ Gemma 3 27B 30.23 - 136.43 - LVLM-Count (Gemma 3 27B as LVLM) 19.70 $\downarrow 10.53$ 110.62 $\downarrow 25.81$ Qwen2 VL 72B AWQ 33.45 - 145.90 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 22.67 $\downarrow 10.78$ 124.73 $\downarrow 21.17$
@@
Table 1: Evaluation on the test set of the FSC-147 dataset. In all tables, the results for base LVLMs and LVLM-Count are reported over three trials. Also, columns marked with $\Delta$ show the performance difference between the LVLM-Count and the base LVLM it uses. Green indicates improvement, while red represents degradation. To see the measured accuracy metrics for this dataset and the subsequent benchmarks, refer to Appendix I. Additionally, MAE analysis across different intervals of ground truth values for FSC-147 is provided in Appendix H.
PASCAL VOC Benchmark. We evaluate the performance of three LVLMs—GPT-4o, Qwen2, and Gemma 3—on this benchmark, both with and without LVLM-Count. As shown in Table 2, LVLM-Count consistently outperforms the base LVLMs.
@@
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ Number of the target segmentation masks 4.03 - 7.47 GPT-4o counting target segmentaion masks 6.60 - 15.09 GPT4o 4.46 - 8.35 - LVLM-Count (GPT4o as LVLM) 3.31 $\downarrow$ 1.15 7.11 $\downarrow$ 1.24 Gemma 3 27B 3.37 - 7.49 - LVLM-Count (Gemma 3 27B as LVLM) 2.87 $\downarrow$ 0.5 5.84 $\downarrow$ 1.65 Qwen2 VL 72B AWQ 4.83 - 8.84 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 4.05 $\downarrow$ 0.78 7.72 $\downarrow$ 1.12
Table 2: Evaluation on the PASCAL VOC counting benchmark.
Emoji-Count. We evaluate the performance of the LVLMs and their performance using LVLM-Count on the Emoji-Count benchmark. The results are shown in Table 3. This is a challenging benchmark, as it requires understanding complex concepts. We observe that taking the number of masks as the final count performs particularly poorly, as for any object of interest in the image, the segmentation stage tends to segment all the objects and cannot distinguish between different icons. Although GPT-4o and Gemma 3 show reasonable performance, the other open-source model does not perform well. Nonetheless, the performance of all three base LVLMs is significantly enhanced by LVLM-Count, especially Qwen2, which performs almost on par with GPT-4o when LVLM-Count is used for it.
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ Number of the target segmentation masks 82.47 - 107.98 - GPT-4o counting target segmentaion masks 107.72 - 162.12 - GPT-4o 22.51 - 35.94 - LVLM-Count (GPT-4o as LVLM) 16.29 $\downarrow 6.22$ 32.47 $\downarrow 3.47$ Gemma 3 27B 21.43 - 24.15 - LVLM-Count (Gemma3 27B as LVLM) 15.76 $\downarrow 5.67$ 21.00 $\downarrow 3.15$ Qwen2 VL 72B AWQ 82.41 - 186.32 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 23.74 $\downarrow 58.67$ 40.04 $\downarrow 143.84$
Table 3: Evaluation on the Emoji-Count benchmark.
Penguin Benchmark. We report the counting performance on the Penguin benchmark in Table 4. We evaluate two variants of LVLM-Count. The primary variant uses GroundingDINO in both the area detection and target segmentation stages. The alternative variant does not use a detection model; instead, SAM is configured to segment all entities in the image. Both variants improve MAE across all three LVLMs, highlighting LVLM-Count’s robustness in scenarios with heavy occlusion and complex backgrounds—conditions that pose significant challenges for area detection and target segmentation. Additional details and a visual example showing the segmentation masks and division paths generated in both variants for a sample from the Penguin benchmark can be found in Appendix D.
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ GPT4o 36.01 - 46.64 - LVLM-Count( Main variant, using GPT-4o ) 26.95 $\downarrow$ 9.06 39.32 $\downarrow$ 7.32 LVLM-Count(SAM-only variant, using GPT-4o) 28.79 $\downarrow$ 7.22 45.71 $\downarrow$ 0.93 Gemma 3 27B 49.87 - 61.06 - LVLM-Count( Main variant, using Gemma 3 ) 36.75 $\downarrow$ 13.12 47.24 $\downarrow$ 13.82 LVLM-Count( SAM-only variant, using Gemma 3 ) 41.97 $\downarrow$ 7.90 54.55 $\downarrow$ 6.51 Qwen2 VL 72B AWQ 44.96 - 66.78 - LVLM-Count( Main variant, using Qwen2 ) 28.67 $\downarrow$ 16.29 43.71 $\downarrow$ 23.07 LVLM-Count(SAM-only, using Qwen2 ) 38.29 $\downarrow$ 6.67 51.47 $\downarrow$ 15.31
Table 4: Evaluation on the Penguin benchmark.
Limitations
In this work, we quantitatively evaluated the visual counting performance of several LVLMs on multiple datasets. More importantly, we introduced a simple yet effective baseline method, LVLM-Count, that enhances the visual counting capabilities of LVLMs across the evaluated benchmarks. However, like any method, LVLM-Count has limitations. One limitation arises in cases where sub-images contain no objects of interest: the LVLM may occasionally predict a non-zero value. This reflects a broader weakness in LVLMs, and addressing it will require targeted improvements to ensure their accurate zero prediction in such scenarios. Another limitation occurs with open-source models during the target counting stage. Ideally, the LVLM’s output for each sub-image should be a numerical value. While top proprietary LVLMs, such as GPT-4o, offer functionalities like JSON schema to enforce structured responses (e.g., ensuring a numerical output), open-source models typically lack such features. For these models, the prompt must include explicit instructions to format the response correctly. For instance, an instruction like "Place the final predicted number inside [[]]" enables the use of regex searches to extract the number for aggregation.
Appendix Table of Contents
[…]
[sections]
[…]
[sections]l1